Velocity v of wave produced in the water depends on the wave length lamda (L) density of water d and acceleration due to gravity g relation betweeen these quantities will be

Velocity v  of wave produced in the water depends on the wave length lamda (L) density of water d and acceleration due to gravity g relation betweeen these quantities will be

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

This is a problem of dimention

let

V = k λadbgc

 

  where k ,a,b,c are constant

Now write the dimension of all the quantities

[ M0 L1 T-1]  = [L1]a[ M1 L-3]b [L1 T-2]c

[ M0 L1 T-1]  = [ Mb L-3b+a+c  T-2c]

compair

b=0

-2c=-1   or c=1/2

and  -3b +a+c =1

 a =1/2

so V =k(λg)1/2

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin